# A Visual Case Study of the Training Dynamics in Neural Networks

### v1

@@ 2.3 Visualization Tools
3 Family of Circuits
In this section, we describe idealized circuits designed to solve the sparse modular addition problem and discuss concrete solutions implemented by transformers.
Figure 1: Idealized circuit capturing the invariants of the problem
@@ 3 Family of Circuits
3.1 Idealized Circuits
  • “Permutation invariance:” $y$ does not depend on permutation of non-spurious tokens, i.e., if $\pi\in\mathfrak{S}_{k}$ is a permutation, then $y=f^{*}((x_{i}))=f^{*}((x_{\pi(1)},\ldots,x_{\pi(k)},x_{k+1},\ldots,x_{L}))$.
[…]
Figure 4 illustrates this ideal model with our visualization. From left to right, we first plot the two token embeddings $E(x)$ for $x\in[p]$ with $p=2$. We then plot the twelve position embeddings $P(t)$. The spurious (resp. non-spurious) positions are represented with squares (resp. circles), they all collapse to a single point. The normalized embeddings $Z(x,t)$ are plotted in the third frame, annotated with $(x,t)$, with the query vector $q$ represented as a red arrow. This arrow points in the direction of the embedding $Z(x,t)$ for $t\in[k]$, allowing the attention mechanism to focus exclusively on non-spurious tokens. Finally, we plot some sequence embeddings, annotated with $(x_{t})$, and the output of the feedforward transform. The feedforward layer is then able to map each cluster to the appropriate output class. Its output is a probability vector over the classes, which we map to a color according to the RGB color wheel.
3.2 Concrete Realizations
In practice, the model weights learned through gradient descent deviate from the idealized model. We observe several variations of the idealized model, which we categorize as follows:
Faulty Attention, Corrected by Value.
In many instances, we find that the attention scores are not fully concentrated on the first five tokens. However, this faulty attention pattern is compensated by the value matrix, which effectively collapses the embeddings of the spurious tokens that are attended to. This adjustment allows the sequence embeddings to remain invariant to the suffix of the sequence. An example of such a configuration is depicted in Figure 4. We also observe examples where the attention focuses solely on non-spurious tokens that are not, for example, zeros, which still allows the sequence embedding to encode for the prefix sums.
Partially Learned Invariants.
Frequently, we observe that the sequence embeddings have not fully learned all the suffix and prefix invariants, resulting in more than six clusters of sequence embeddings. Specifically, Figure 4 illustrates a scenario where the sequence embeddings are not invariant to the value $x_{6}$, as evidenced by the positions of the blue squares $(0,6)$ and $(1,6)$ on the plot. They also lack invariance to permutations of the token in the first or fifth positions with another of the non-spurious tokens. This results in a sequence embedding that presents more clusters than the idealized model, leading to a greater number of connected decision regions in the feedforward layer.
Fuzzy Constructions.
Occasionally, we encounter fuzzy constructions where the sequence embeddings are clustered according to unconventional patterns that nonetheless generalize to unseen data. Such a construction is presented at the bottom of Figure 4.
To conclude, we observe numerous variations from the idealized construction. The networks discover a variety of weight configurations to effectively solve our task. While these configurations do not exhibit particularly striking geometric structures, they often capture the invariants of our problem.
4 Training Insights
This section discusses several training insights we obtained from our sandbox. It focuses on training dynamics, the transferability of circuits, and loss spikes.
4.1 Training dynamics
Our toolbox enables us to precisely track the training dynamics of the model. We notably observe that the loss curves typically present drops, corresponding to the learning of different parts of the network, hierarchically from the first to the last layers. They are illustrated in Figure 7.
Figure 6: Ablation studies regarding test accuracy as a function of batch size, hidden neurons, learning rates, and MLP learning rates discount factor for a single run with 1000 epochs. We ensured consistency in initial weights and batch designs when changing hyperparameters.
Figure 7: This study, akin to Figure 7, yet averaged over $100$ runs, highlights some regularity in the effect of hyperparameters on the resulting test accuracy. The solid color indicates standard deviations.
First loss drop: learning of the sequence embeddings.
The first loss drop coincides with the learning of the sequence embeddings. It really corresponds to a phase change in the dynamics of the weights. Before the first phase change, the weights seem to wander as if trapped in a saddle point, waiting for a clear signal to escape it. At one point, they all move quite rapidly to create a relatively definitive structure for the sequence embeddings. Interestingly, we also notice that when changing the training hyperparameters, the time for this phase change to occur can vary quite a lot, reflecting the highly unpredictable time needed to escape from the saddle point. Figure 7 shows the high variability of the test accuracy after 1000 epochs when slightly varying one hyperparameter (Figure 7 shows more regularity when averaging over the runs). The exit from a saddle point can also be understood via the gradient norms as illustrated in Figures 16 and 16 of Appendix A.2. Our findings resonate with some theoretical arguments found in the deep learning theory literature (Dauphin et al., 2014; Du et al., 2017; Chi et al., 2019), results that we hope our toolbox could help theorists strengthen.
Second loss drop: fitting of the MLP.
The second loss drop is due to the learning of the feedforward network. This change is about fitting the MLP weights to assign the correct classes to the different clusters created during the learning of the sequence embeddings. Interestingly, this second loss drop does not correspond to a clear phase change in the dynamics of the weights. The MLP weights seem to evolve at a continuous speed, although the corresponding decision frontiers change relatively strongly. We notice that this second loss drop appears soon after the first one, if not simultaneously. Once again, we hope that theorists could use our visual toolbox to shed more light on this hierarchical learning of the weights and strengthen previous results on the matter (see, e.g., Abbe et al., 2021).
The influence of initialization.
As we have seen in the previous subsection, the final configuration can vary quite a lot from one run to another. From our manual inspection, the final configuration seems to be highly correlated with the initial weight configurations. We notably see a strong similarity between the attention patterns at the start and at the end of the training, as illustrated by Figure 8. We also notice that modifying the training hyperparameters (batch size, learning rates, etc.) without changing the initial weights does not significantly alter the final weight configurations.
(a) Pretraining with $p=2$.
(b) Finetuning with $p=3$.
Figure 8: Influence of the initialization. Visualization of attention maps: for each sequence on the left, we plot the corresponding attention pattern at the start and the end of training. The final attention maps are correlated with the original one, illustrating that final variations of the original circuit depend on the original weight configuration.
4.2 Transfer Experiments

### v2

@@ 2.3 Visual Sandbox
3 Clustering Head: A Circuit Perspective
In this section, we adopt a circuit perspective and describe an efficient pathway for solving the sparse modular addition problem. In particular, we describe an idealized “clustering head” circuit, as well as variants observed in practice.
3.1 Idealized Circuit
  • “Permutation invariance:” $y$ does not depend on permutation of non-spurious tokens, i.e. for $\pi\in\mathfrak{S}_{k}$, $f^{*}((x_{i}))=f^{*}((x_{\pi(1)},\ldots,x_{\pi(k)},x_{k+1},\ldots,x_{L}))$.
These two sets of invariants can be easily enforced by the embedding layer. Any architecture where the position embeddings satisfy $p_{t}=p_{1}$ for $t\in[k]$, will be permutation invariant, meaning that its output will be invariant to permutations of the non-spurious tokens. Similarly, suffix invariance can be enforced by ensuring that the query vector primarily aligns with $Z(x,t)$ for $t\in[k]$, allowing the sequence embedding $\xi$ to be invariant to the sequence suffix. Such a construction would yield (${k+p-1\choose k}$ clusters of sequence embeddings, [^2] This number corresponds to the number of ways to split $k$ into $p$ buckets, which is also the number of stars and bars configurations with $k$ stars and $p-1$ bars. which the feedforward layer could scatter into as many decision regions to map each sequence embedding $\xi$ to the correct output class $y\in[p]$. The pathway that realizes such operations is called a Clustering Head.
Figure 1 illustrates this ideal model using our visual sandbox. From left to right, we first plot the two token embeddings $E(x)$ for $x\in[p]$ with $p=2$. We then plot the twelve position embeddings $P(t)$. The spurious (resp. non-spurious) positions are represented with squares (resp. circles), they all collapse to a single point. The normalized embeddings $Z(x,t)$ are plotted in the third frame, annotated with $(x,t)$, with the query vector $q$ represented as a red arrow. This arrow points in the direction of the embedding $Z(x,t)$ for $t\in[k]$, allowing the attention mechanism to focus exclusively on non-spurious tokens. Finally, we plot the sequence embeddings $\xi$ of some $(x_{t})$, and the output of the feedforward transform. The embeddings are clustered and respect the invariance, and the feedforward layer can map each cluster to the appropriate output class. Its output is a probability vector over the classes, which we map to a color according to the RGB color wheel.
3.2 Concrete Realizations
We expect clustering heads to naturally appear during training given that they are an efficient and simple way to solve the task. In practice, the sequence embeddings $\xi$ learned through gradient descent can be grouped in similar but different ways, which leads to several variations of the idealized head. These variations are categorized below.
Faulty attention corrected by value.
In many instances, we observe that the attention scores are not fully concentrated on the first five tokens but are compensated by the value matrix, which effectively collapses the embeddings of the spurious tokens that are attended to. This allows the sequence embeddings to remain invariant to the suffix of the sequence as can be seen in Fig. 13.
Partially learned invariants.
Frequently, the sequence embeddings have not fully learned all the suffix and prefix invariants, resulting in more than six clusters of sequence embeddings [^3] This corresponds to (${k+p-1\choose k}$ for $p=2,k=5$.. Specifically, Figure 13 illustrates a scenario where the sequence embeddings are not invariant to the value $x_{6}$, as evidenced by the positions of the blue squares $(0,6)$ and $(1,6)$ on the plot. They also lack invariance to permutations of the token in the first or fifth positions with another of the non-spurious tokens. This results in a sequence embedding that presents more clusters than the idealized model, leading to a greater number of connected decision regions in the feedforward layer.
Fuzzy constructions.
Occasionally, we encounter fuzzy constructions where the sequence embeddings are clustered according to unconventional patterns that nonetheless generalize to unseen data. Such a construction is presented in Figure 13.
4 Learning Clustering Heads
This section investigates the training dynamics of solving the sparse modular addition problem. We begin by making empirical observations: we observe that transformers typically learn clustering heads through a two-stage learning process and that loss spikes can occur during this process. We then discuss theoretical insights that can be derived without any model simplification assumptions.
4.1 Two-Phase, One Loss: Escaping Saddle Points
A first experimental observation is the typical two-stage learning process that transformers undergo in our setting: first clustering sequence embeddings, before fitting a classifier on top. In particular, the loss curves typically present two drops, corresponding to the learning of different parts of the network, hierarchically from the first to the last parts of the network. This is illustrated in Figure 2. Our findings resonate with theoretical arguments in the deep learning theory literature on saddle point (Dauphin et al., 2014; Du et al., 2017; Chi et al., 2019a) and hierarchical learning (Abbe et al., 2021).
First loss drop: learning of the sequence embeddings.
The first loss drop coincides with the learning of the sequence embeddings. It corresponds to a phase change in the dynamics of the weights. Before the first phase change, the weights seem to wander as if trapped in a saddle point, waiting for a clear signal to escape it. At one point, they all move quite rapidly to create a relatively definitive structure for the sequence embeddings. Interestingly, we also notice that the time it takes for this phase change to occur when changing the training hyperparameters can vary quite a lot, reflecting the highly unpredictable time needed to escape from the saddle point. This is illustrated in Section A.4 where Figure 16(a) shows the high variability of the test accuracy after 1000 epochs when slightly varying one hyperparameter (Figure 16(b) shows the same phenomenon with more regularity when averaging over the runs). The exit from a saddle point can also be understood via the gradient norms, as discussed at the end of this section.
Figure 3: Connection to saddle points. From left to right: Evolution of train and test losses, the corresponding accuracies, the evolution of gradient norms for each layer in log-scale, and the similar evolution in linear scale in full-batch. We see that the learning phases of Section 4.1 appear in tandem with high gradient norms. This can be seen in the last subfigure where the three pics correspond to the loss drops and their corresponding plateaus.
Figure 4: High-curvature feed-forward. Loss spikes (both left) resulting from a small change from one iteration to another in sequence embeddings that are close to the decision boundaries of the subsequent feedforward layer (both right).
Second loss drop: fitting of the MLP.
The second loss drop is due to the learning of the feedforward network. This change is about fitting the MLP weights to assign the correct classes to the different clusters created during the learning of the sequence embeddings. Interestingly, this second loss drop does not correspond to a clear phase change in the dynamics of the weights. The MLP weights seem to evolve at a continuous speed, although the corresponding decision frontiers change relatively strongly. We notice that this second loss drop appears soon after the first one, if not simultaneously. When the model’s training stagnates, the loss plateaus for several epochs before decreasing again. This behavior may be understood by the optimizer reaching a saddle point in the loss function w.r.t. the model’s parameter, that is the gradient of the loss is (almost) zero, but the optimization has not yet reached a local minimum. The connection between gradient norms and learning phases is salient in Fig. 3 where we can see that loss drops occur in tandem with high gradient norms for each layer.
4.2 Occurence of Loss Spikes
Better understanding why loss spikes occur is of great interest as they can be detrimental to the training stability, particularly with very large-scale transformers such as large language models (Chowdhery et al., 2023). While some know-how exists for mitigating the impact of peak losses, it is often experimentally motivated and requires a lot of computational resources to achieve operational success (Grattafiori et al., 2024). One interesting aspect of our sandbox is that it generates loss spikes that we can study quite precisely. Our visual inspection showcases two aspects leading to loss spikes: the high curvature of the RMS normalization layer near the origin as well as the high curvature of multi-layer perceptrons with heavy weights, or with numerous small correlated weights.
High-curvature normalization.
Figure 5 illustrates that a small modification to an element can result in a disproportionately large change in its normalized version. In theory, gradient points towards directions that would reduce the training loss. However, considering a large step size in these directions could be counterproductive. This is especially true for functions with high curvature, such as the normalization layer $f(x)=x/\left\|x\right\|$ near the origin. At any point $x_{t}$, the gradient descent update rule suggests that one can update $x_{t+1}$ as $x_{t}-\eta_{t}u_{t}$ without changing $f(x_{t})$, where $u_{t}=x_{t}/\|x_{t}\|$ and $\eta_{t}$ is the learning rate. This holds true only if the learning is small enough, $\eta_{t}<\|x_{t}\|$. When $x_{t}$ is close to zero, ensuring $\eta_{t}<\|x_{t}\|$ becomes challenging, particularly if the step size $\eta_{t}$ was predetermined by some scheduler. This behavior is related to the "edge-of-stability" phenomenon highlighted by Cohen et al. (2022), further theoretical insights being provided by Cabannes et al. (2024c). Interestingly, this analysis suggests removing some loss spikes by smoothing out the normalization layer. For example, consider using $f(x)=\sigma(\left\|x\right\|)x/\left\|x\right\|$ where $\sigma$ is a smooth function with $\sigma(0)=0$ and $\sigma([1,\infty))=\{1\}$. This demonstrates the usefulness of our visual sandbox in gaining insights and building intuition, which can then be validated on a larger scale in subsequent works.
Figure 5: High-curvature normalization. Loss spikes are linked to the high curvature of internal network functions. A small update to an element $x$ can result in a substantial change to its normalized version $x/\left\|x\right\|$, significantly altering the network’s subsequent behavior.
High-curvature feed-forward.
Connection to gradient norms.
Finally, we observe in Fig. 7 that the loss spikes appear in tandem with high gradient norms, indicating that when a too-large step size deviates the model from its current small loss region, it is taken back to where it was with large updates. We note that we observe similar patterns in larger models with embeddings size $d>2$ (see Appendix B). It has been shown in the literature that studying the gradient and the hessian of the loss could provide valuable insights on the neural network optimization both from a theoretical and experimental point of view (Ilbert et al., 2024; Foret et al., 2021; Zhang et al., 2024b). We believe using our visual sandbox could help in future work on such an investigation.
Figure 7: Loss spikes and gradient norms. This study, akin to Fig. 3 in mini-batch, makes the connection between gradient norm and loss spikes salient as loss spikes occur in tandem with spikes in gradient norms.
4.3 Further Observations
Our sandbox allows us to make other interesting observations, such as the impact of initialization on the type of final circuits, the importance of active neurons, and the usefulness of curriculum learning to induce the correct invariance.
Impact of the initialization.
We observe in Fig. 8 the strong similarity between the attention patterns at the start and at the end of the training, indicating that initialization plays a crucial role in the concrete realizations of clustering heads. We note that the transformer learns to focus on the $1$ among the non-spurious tokens, before counting them and deducing the number of 00 to make its final prediction.
Figure 8: Impact of the initialization. For each sequence (left), we plot the corresponding attention pattern at the start (middle) and the end of training (right). The darker, the higher the attention weight. The final attention maps are correlated with the original one, illustrating that variations of the idealized circuit depend on the original configuration. We see that the transformer learns to focus solely on the first $k=5$ input tokens, which are the ones defining the output $y$.
Figure 9: This study is akin to Fig. 8 when the model is first pre-trained with $p=2$ and then finetuned with $p=3$. This illustrates that the final variations of the original circuit depend on the original weight configuration. We notice that the model learns better to focus on the non-spurious tokens.
Figure 10: Connection between performance and sparsity. Since the activation is the GELU that does not map all the negative entries to 00 (contrary to the ReLU), we define the activation sparsity as the percentage of entries with values lower than $\varepsilon>0$. We display the evolution of the activation sparsity of $20$ trained models with $\varepsilon\in[10^{-5},10^{2}]$. Left: Successful models (i.e., with test accuracy above $0.9$) in blue have less sparse activation than failed models in red. Right: The color indicates the models’ test accuracy (the lighter, the better). The performance increases as the activation sparsity decreases.
Transferability of circuits.
Among the lessons of training very large models is that some sources of data may facilitate the learning of certain skills (Touvron et al., 2023b; Llama Team, 2024; Cabannes et al., 2024a). In Section A.3, we illustrate this phenomenon in the learning of clustering heads. In particular, we show that training first a model on a simpler task (e.g., $p=2$) helps it better solve a more challenging one (e.g., $p=3$) as can be seen in Fig. 9 along with the impact of initialization. Although in two stages, our experiments might indicate that the transferability of circuits can explain the importance of data curation in more practical settings.
Connectivity of neurons.
As explained in Section 3.1, clustering heads capture the invariance of the sparse modular addition problem. Motivated by recent works studying the sparsity of transformers’ activations (Li et al., 2023a; Mirzadeh et al., 2024), we show in Fig. 10 that the models that successfully solve the task have denser activation. This might indicate that many neurons are necessary to correctly understand the invariance, hence approximating the idealized circuit well (more details in Section A.2).
Generalization to bigger embeddings.
While the previous observations were enabled by our toolbox, which precisely tracks the training dynamics when embeddings reside in ${\mathbb{R}}^{2}$, our findings are confirmed when considering larger embedding dimensions in Appendix B. In particular, we discuss the two-phase learning in Section B.1, the loss spikes in Section B.2, and the activation sparsity in Section B.3. We also note the coherence of our results with existing literature in more practical settings (Li et al., 2023b; Mirzadeh et al., 2024; Zhang et al., 2024b; Ilbert et al., 2024; Cabannes et al., 2024c).
4.4 Theoretical Insights
In this section, we provide theoretical insights on the learning of clustering heads. In practice, our transformer is trained via gradient descent which converges to the stationary points of $\mathcal{L}$, i.e., solutions of $\nabla\mathcal{L}=0$. The closed-form gradients are given in Eq. 5 and these stationary points are characterized exactly in C.2. They notably include the clustering heads construction, and all their variants which allow them to reach the loss minimum ${\mathcal{L}}^{*}=H(Y\mid X)$ which is the conditional entropy of $Y$ given $X$. To better understand the two-staged dynamic observed in Section 4.1, we investigate in Section 4.4 how the gradient magnitude is controlled along the training. The formal statement is deferred to Section C.4. {boxprop}[Gradient upper-bound, informal] Assuming bounded token embedding, the gradient norm is upper-bounded by
$\lVert\nabla\mathcal{L}(\theta)\rVert\leq\mathcal{O}\left(\tilde{B}\cdot\sqrt{\mathscr{E}}\right),$
where $\mathscr{E}=\mathbb{E}_{x}\left[d_{\mathrm{TV}}\left(\hat{p}_{\theta}(\cdot|x),p(\cdot|x)\right)\right]$ captures the ability of the transformer to solve the task and $\tilde{B}$ depends solely on the spectral norm of the model’s parameters. Since our model does not include any theoretical simplification to the transformer architecture, the proofs are somewhat involved. Section 4.4 is useful to understand the loss profiles identified in Section 4.1. Notably, we see that the term $\mathscr{E}$ captures the ability of the model to correctly solve the task. In the early iterations, the model behaves poorly which imposes a weak control on the gradient norm. This allows for large gradient norms and hence for bigger updates to the weights. The loss can vary rapidly as can be seen by the steep descent of Fig. 2. During training, the model improves as can be seen by the increase in accuracy (both train and test). This leads to a smaller $\mathscr{E}$ and strengthens the control on the norm. In parallel, the gradient updates are of smaller magnitude as well as the loss variations. This implies a less steep descent in the loss as observed in Fig. 2.
5 Discussion
This paper aims to advance our understanding of the transformer’s training dynamics through a detailed study of a small-scale model. We introduced a visual sandbox that provides comprehensive visualizations of each layer throughout the training process. It allowed us to reveal a type of circuit, called "clustering heads," which learns the invariants of the problem. We analyzed the learning dynamics of these circuits, showcasing two-stage learning, and loss spikes caused by high curvature in the model’s internal functions. In addition, we studied the effect of initialization and curriculum learning. Our theoretical insights highlighted that clustering heads can indeed be learned by gradient descent and provided an understanding of the two-stage learning observed in practice.
Future work.
Future work will aim to apply these insights to larger and more complex models, assess the scalability of observed phenomena, and enhance our visualization tools for higher-dimensional models.
Impact Statement
This paper aims to understand the complex dynamics at play in transformers. The insights gained from our research stream could help design more efficient and robust models and optimization schemes, ultimately aiding in reducing the computational cost and environmental impact of training large models.
Acknowledgements
We thank Elvis Dohmatob, Eshaan Nichani, Giuseppe Paolo, Faniriana Rakoto Endor, and Ievgen Redko for the fruitful discussions that led to this work.
Author Contributions
Ambroise worked on the codebase, the experiments, the proofs, and the writing. Wassim worked on the codebase, the experiments, and the writing. Vivien proposed the initial idea, worked on the codebase, and supervised the project. Each author was essential to this work.
Appendix
Roadmap.
We provide additional experiments in Appendix A. We go beyond our controlled setting and present results with higher-dimensional embeddings in Appendix B. In Appendix C, we discuss in detail the theoretical insights on the learning dynamics. Finally, the proofs of our theoretical results are given in Appendix D. The corresponding table of contents is presented below.
Appendix A Experimental Study of Clustering Heads
In this section, we present additional results related to the study of the circuits implemented by the transformer. Notably, we investigate the concrete realizations to solve the task, the sparsity of the model’s activation along the training, and the transferability of circuits.
Implementation details.
We open-source the code at https://github.com/facebookresearch/pal. The implementation details are given in Section 2 along with the default configuration below. In our experiments, we trained our networks using $n=2048=2^{11}$ data points, which were sampled uniformly with replacement from the $p^{L}$ possible sentences. We utilized the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ (Kingma & Ba, 2017), and we initialized the network weights using the default schemes provided by PyTorch (Paszke et al., 2019).
A.1 Concrete Realizations
In practice, the model weights learned through gradient descent deviate from the idealized model. In this section, we visualize the variations of the idealized circuit described in Section 3.2 and categorized as follows:
  1. 1.
    Faulty attention corrected by value: the attention is compensated by the value matrix to capture well the invariance,
  2. 2.
    Partially learned invariants: the sentence embeddings present more clusters than the idealized model,
  3. 3.
    Fuzzy constructions: the clustering is unconventional but the model still generalizes well.
@@ A.1 Concrete Realizations
A.2 Connectivity of Neurons
Sparsity is a phenomenon of interest in many fields such as signal processing, neuroscience, and machine learning (Mairal et al., 2009; Chen et al., 1998; Barth & Poulet, 2012). Recent studies focused on the sparsity in deep neural network activations. In particular, Li et al. (2023a) showed that trained transformers have sparse activations and concluded that it was caused by the training dynamics rather than by a compact representation of the training data as commonly thought in computer vision and NLP. Mirzadeh et al. (2024) observed a similar phenomenon and showed how to leverage sparsity to reduce the inference cost of large language models. Inspired by this line of work, we study the activation sparsity of our model from a performance viewpoint. It should be noted that those works study deep transformers and identified that the sparsity increases with the depth while we only consider a one-layer transformer. In our setting, the activation sparsity can help us better understand how neurons connect to implement clustering heads.
Successful models have dense activations.
Following the framework from Li et al. (2023a), we recall that the activation sparsity corresponds to the percentage of non-zero entries of the feed-forward activation map. Without loss of generality, the feed-forward block is an MLP with weights $W_{1},W_{2}$ and a non-linear activation $\sigma$ that outputs for any input $x$ a vector $z=W_{1}\sigma(W_{2}x)$. Formally, the activation sparsity is the percentage of non-zero entries in $\sigma(W_{2}x)$ and takes values in $[0,1]$. In the classical setting with ReLU activation (Fukushima, 1969), this is equivalent to computing the percentage of non-negative neurons before the activation. However, some activations do not have non-negative outputs. This is the case of the SiLU (Elfwing et al., 2018), used in Llama models (Touvron et al., 2023a), and of the GeLU Hendrycks & Gimpel (2023) used in Falcon (Almazrouei et al., 2023), PaLM (Chowdhery et al., 2024), and in our transformer implementation. Instead of replacing such activations with a ReLU (Mirzadeh et al., 2024; Li et al., 2023a), we compute a smoothed sparsity. This is akin to using the $\ell_{1}$-norm, respectively the nuclear norm, instead of the $\ell_{0}$ quasi-norm, respectively the rank (Ilbert et al., 2024; Gribonval & Nielsen, 2003). as the percentage of entries with an absolute value lower than $\varepsilon>0$. A sparsity of $1$ means that all entries are $\varepsilon$-close to 00 (i.e., sparse activations) and a sparsity of 00 means that all entries are at least $\varepsilon$-away from 00 (i.e., dense activations).
Evolution of the sparsity with performance.
To better understand the impact of sparsity, we train $20$ independent models and display in Fig. 10 their sparsity after training for $\varepsilon\in[10^{-5},10^{2}]$ (the range is chosen such that the sparsity reaches its extremal values 00 and $1$). Given the task’s difficulty, achieving an accuracy above $0.9$ is a success; otherwise, it is a failure. On the left, we plot successful models in blue and failed ones in red. We observe a striking separation between successful and failed training. In the permissible range $[10^{-3},10]$, successful models tend to have less sparse activation than failed ones. To further study this phenomenon, we plot in the right subplot of Fig. 10 the evolution of the sparsity with $\varepsilon$, and here, the color indicates the models’ test accuracy (the lighter the color, the better the model). We can see that the sparsity decreases as the performance increases. This explains the sharp transition between failure and success in the left subplot. These experiments seem to indicate that, contrary to images and textual data (Li et al., 2023a; Mirzadeh et al., 2024), the sparse modular addition problem needs the involvement of many neurons during inference, and hence requires non-sparse activations.
A.3 Transferability of Circuits